OpenStudy (anonymous):

\[4b^{2} + 8b = 0\] Please show me the steps to solve

OpenStudy (anonymous):

4b(b+2)=0 so 4b=0 or b+2=0 b=0 or b= -2

OpenStudy (anonymous):

can you further explain why and how? like walk me through step by step.

OpenStudy (anonymous):

is it that we break down 4b^2 into ( 4b ) ( b ) and we devide 8b by 4b and get 2b? but on which side of the parentheses does the 8 go, i really need it written out so i can follow step by step.

OpenStudy (anonymous):

no 4b is a factor of 4b^2 and 8b. therefore it is factored out

OpenStudy (anonymous):

4b(b+2)=0 is factoring when you multiply 4b times b + 2, it will = 4b squared + 8b which was your original problem. It is like simplifying that is required to get your answer.

OpenStudy (anonymous):

after you factor the problem out which is 4b(b+2)=0 you have to separate the two parts of your factoring. 4b is one part that is multiplied by the other part which is b+2. separately set each part equal to zero. 4b=0 and b+2=0 and solve for b

OpenStudy (anonymous):

From my point of view I am trying to see how i would write out problems similar to these. the steps, when do i put in parentheses , how do i know when to put +or - if you write it out step by step i could just copy and see it.

OpenStudy (anonymous):

Sorry, if i am being difficult I just want to have an understanding.

OpenStudy (anonymous):

it's fine, understanding it is much better than just being given an answer. i'm drawing it out

OpenStudy (anonymous):

Thanks, I already have the answer it just bugs me if i can't find the way.

OpenStudy (anonymous):

It's going to take a little while because it's hard to write/ draw on the computer, is that okay?

OpenStudy (anonymous):

yeah, that is fine . In the mean time I am working it out xD

OpenStudy (anonymous):

I think i have to subtract 4 b both sides, other wise if i foil it back i don't get -2

OpenStudy (anonymous):

you don't subtract the 4 but you do foil when checking your factoring (factoring is guess and check pretty much)... in 4b(b+2)=0 you foil to figure out if it's the correct factor. 4b*b=4b^2 and 4b*2=8b. The plus sign is there because you can't add b^2 to b which means it's already simplified, so when you foil that factor, you get what you started with which was 4b^2+8b=0. Because you get what you started with from foiling, that is how you know your factoring is right. Since 4b(b+2)=4b^2+8b when foiled and 4b^2+8b=0, 4b(b+2)=0. Since you get what you started with when multiplying 4b to b+2, you have to separate those two parts and make each part equal zero. Then you have 4b=0 and b+2=0, so you will have two answers. You get the first answer from 4b=0 by isolating b by dividing both sides by 4. dividing 4b by 4, you get b and by dividing 0 by 4, you get 0. so the answer to that would be b=0. Your second answer comes from b+2=0 by isolating b again by subtracting 2 from both sides. b+2-2 is just b and 0-2 is -2. so your second answer would be -2. Does that help a little more?

OpenStudy (anonymous):

I'll try to draw it too

OpenStudy (anonymous):

Thanks! this helps a lot, i mean a lot!

OpenStudy (anonymous):

your welcome :D

OpenStudy (anonymous):